LIFETIMES OF THE Ã STATES OF C_{3}

YI-JEN WANG, CHIAO-WEI CHEN, LIUZHU ZHOU, ANTHONY J. MERER, \underline{YEN-CHU HSU}, Institute of Atomic and Molecular Sciences, Academia Sinica, P. O. Box 23-166, Taipei 10617, Taiwan, R. O. C..

Fluorescence lifetimes have been measured for 66 Π_{u} and 16 Σ_{u}^{+} vibronic levels of the Ã^{1}Π_{u} state of C_{3} under nearly collision-free conditions. The vibrational levels studied were v_{1}=0−5, v_{2}=0−22, and v_{3}=0−4, covering the energy range up to 6000 cm^{−1}, and for most of them only the R(0) line was investigated. The measurements are estimated to be accurate to better than ±3 ns. General features of the lifetimes are as follows: Pure bending levels with Π_{u}-symmetry have shorter lifetimes (as short as 190 ns). With increasing excitation of the stretching vibrations, the lifetimes increase. Bending excitation generally reduces the lifetimes, though exceptions occur for v_{1}=1. Levels with v_{1}+v_{3} ≥ 3 or with vibrational energy greater than 4000 cm^{−1} have lifetimes longer than 300 ns; some of these levels even have a long-lived component with a lifetime of up to several microseconds. This implies that the predicted dark states, (B̃^{1}Σ_{u}^{−} and B̃′ 1Δ_{u}), lie at least 4000 cm^{−1} above the 0 0 0 level of the Ã state. It is interesting to note that most of the Σ_{u}^{+}-symmetry levels lying above 4000 cm^{−1} also have lifetimes longer than 300 ns.